Assume all sales are one-time credit sales with a probability of collection of 96 percent. The variable cost per unit is $1.67, the sales price per unit is $4.99, and the monthly interest rate is 1+x percent. What is the NPV of a credit sale of one extra unit? Let x=1

A) $3.073

B) $3.026

C) $2.981

D) $2.936

E) $2.892
Answer #1

Answer:- Option (B): $3.026

Explanation:-

Net Present Value = { (Probability of collection) x [ Sales price per unit / ( 1+ Discount rate)] } - Variable cost per unit

= [(1-.04) x ($4.99/1.02)] - $1.67

= [(0.96) x ($4.99/1.02)] - $1.67

= $4.696470588 - $1.67

= $3.026470588

OR

= $3.026

[ Note: Discount Rate = monthly interest rate of (1+x) percent = (1+1)% = 2% ]